OpenStudy (goldrush18):

simplify (2a^-1)^2 /(2a^-1)^-2

OpenStudy (anonymous):

oops that is wrong. it is \[(2a^{-1})^4=2^4a^{-4}=\frac{16}{a^4}\]

OpenStudy (goldrush18):

thanks i got that too when i tried it

OpenStudy (goldrush18):

i have more that needs to be corrected if im wrong